OpenStudy (anonymous):

Find the solution of five times the square root of the quantity of x plus 7 equals negative 10 and determine if it is an extraneous solution. x = −3; extraneous x = −3; not extraneous x = 11; extraneous x = 11; not extraneous

OpenStudy (anonymous):

Is this the equation? \[\Large 5\sqrt{x+7}=-10\]

OpenStudy (anonymous):

im not sure lol i hate math so much.. but probably

OpenStudy (anonymous):

Where is your question? Can you show me a picture or a screenshot?

OpenStudy (anonymous):

yeah thats the problem my bad i couldnt find it

OpenStudy (anonymous):

@doulikepiecauseidont thats the problem

OpenStudy (anonymous):

First step to this is to isolate the square root so you divide both sides by 5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what did you get it to =?

OpenStudy (anonymous):

x+7 = -10/5

OpenStudy (anonymous):

well sqrt of x+7= -10/5 but yes, -2=-10/5, so the next step is to square both sides to get rid of the square root, what is (-2)^2?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

what is 2 *2? that is the same as 2 squared

OpenStudy (anonymous):

1.4